# Monitored long-range interacting systems: spin-wave theory for quantum trajectories

### v1

@@ Abstract
We introduce a stochastic spin-wave theory tailored to describe quantum trajectories in continuously monitored long-range interacting spin systems. Our method, based on the bosonization of spin-wave excitations on top of a strong collective polarization, enables the efficient simulation of large-scale interacting spins, offering insights into nonlinear features of the dynamics such as entanglement and trajectory correlations. We showcase the versatility of our framework by exploring an entanglement phase transition in a monitored spin system with power-law interactions and dwelling on how our method mitigates the experimental challenges of post-selection in detecting monitored quantum phases.
@@ I Introduction
Long-range interacting quantum many-body systems have recently been the focus of intense theoretical and experimental activity [1, 2]. Sufficiently non-local interactions between the elementary constituents of a many-body system lead to exotic non-equilibrium phenomena, with compelling signatures in the properties of quantum correlations and entanglement [3, 4, 5, 6, 7, 8], relaxation dynamics [9, 10, 11], quantum information scrambling [12, 13, 14], and ergodicity-breaking properties [15, 16, 17].
Along with their importance in statistical mechanics, long-range quantum systems are central to the rising field of quantum technologies and simulators. Experimental platforms such as trapped ions [18], Bose-Einstein condensates in cavities [19], dipolar [20], polar [21], and Rydberg atoms [22], or driven ultra-cold atomic gases [23] showcase frameworks where the system interactions scale as a power law of the distance. Depending on the experimental platform, the exponent governing the range of the interaction varies and, in some cases, can be tuned, allowing us to explore several different regimes.
@@
The dynamics of long-range systems has been investigated in the opposite limits of unitary and dissipative (Lindbladian) evolutions. In this work, we would like to take a step forward by studying the in-between framework of monitored dynamics. Here, the system evolution is interspersed with quantum measurements, whose outcomes are stochastic. As a result, the system is described by a quantum trajectory conditional to the measurement registry. Averaging the state over the trajectory ensemble recasts a dissipative dynamics, therefore presenting a convenient framework for the study of physical observables, i.e. hermitian operators, in Lindblad dynamics [24, 25, 26, 27, 28]. It was more recently realized, in the study of many-body systems, that the quantum trajectory ensemble contains richer information than the mean state, showcased by the several collective phenomena encoded in beyond-average statistical features. The cornerstone examples are monitored quantum phases and measurement-induced phase transitions [29, 30, 31, 32, 33], characterized by non-linear functions of the trajectories, such as entanglement probes or trajectory correlation functions. Intensive work gathered insights on the monitored phases in local quantum circuits [34, 35, 36, 37, 38, 39], dual-unitary circuits [40, 41, 42], monitored non-interacting [43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55] and interacting [56, 57, 58, 59, 60] systems, interpolating between strong and weak measurements [61, 62], and presenting a comprehensive understanding in terms of quantum error correction and learning capabilities [63, 64, 65, 66, 67, 68, 69, 70, 71].
@@
Experimentally detecting monitored phases, on the other hand, is challenging, and so far limited to few pioneering works on trapped-ions [72] and superconducting platforms [73, 74]. This limitation has a fundamental origin and is known as the post-selection problem. Observing beyond average moments of the quantum trajectories requires that one should reproduce the same sequence of measurement outcomes multiple times to collect enough statistics and obtain trajectory expectation values. This task is formidable, as the probability of reproducing the same trajectory is exponentially small in system size and time scale.
@@
The post-selection barrier is in general ineludible for the study of generic many-body monitored systems, and the quest for methods that can mitigate it is an active research line. For instance, the presence of feedback dynamics may be designed to imprint the measurement-induced transition into the density matrix [75, 76], albeit in general leads to separate types of phase transitions [77, 78, 79, 80, 81, 82, 83, 84]. A complementary path is pursuable for systems that can be classically simulated. The approach combines the outcome of the quantum evolution and the quantum measurement registry, with the classical post-processing fed by the latter. When the post-processing is efficient and faithful, it allows to elude the post-selection [64, 85, 86, 87, 88, 89, 90]. There are special cases in which the system itself is designed to be immune to post-selection. In [91] some of us showed that there is a class of infinite-range spin systems where monitored many-body dynamics can be efficiently realized with a post-selection overhead scaling, at most, as a power of the system size. In Ref. [91], it was argued that this fortunate case was not specific to that model, describing an atomic ensembles driven by a laser field and in the presence of a collective decay, but it applies to a broad class of monitored systems with an underlying semi-classical dynamics.
@@
Few works have addressed monitored dynamics in long-range interacting systems. Refs. [92, 93] studied Clifford circuits with two-qubit gates entangling distant sites with a probability of decaying as a power-law of their distance. The Floquet dynamics of interacting spin systems with measurements and feedback was considered in a model with entanglement and dissipative phase transitions [94], while monitored long-range free fermions have been studied in Refs. [95, 96]. All these results demonstrate that long-range couplings strongly affect the monitored phases and their transitions: they are relevant in the renormalization group sense.
Here we will show how one can study entanglement dynamics as well as phase transitions in long-range spin systems, by designing a stochastic spin-wave expansion along quantum trajectories. This method will be instrumental in showing how to avoid post-selection in long-range systems. Moreover, the method has a much wider spectrum of applications since the stochastic spin-wave quantum trajectories provide a much more accurate solution to the Lindblad master equation as compared to deterministic spin-wave theories [97].
While the many-body problem renders the exact solution of such systems formidable, long-range interactions typically lead to an underlying semi-classical dynamics that alleviates this difficulty. Consider a system with $N$ fully connected, i.e. infinite-range, spin-$1/2$ particles with permutation invariance. The entire system then behaves as a single collective spin $S=N/2$. For large $N$, the collective quantum spin operators, when re-scaled by $S$, admit vanishingly small ($\sim 1/S$) commutators and hence can be well approximated by a classical continuous spin vector [2]. This is essentially the mean-field approximation, which is exact in the $N\to\infty$ limit. On top of this classical spin, quantum excitations and fluctuations can then be treated perturbatively. This argument can be extended to long-range models without permutation symmetry, which, by construction, remain close to their infinite-range counterparts, with non-collective excitations decorating the classical spin. These observations lead to a class of approximation methods known as the spin-wave theory. Historically, the spin-wave theory was first introduced by Bloch [98] in 1932 for ferromagnetic spin systems in equilibrium. An equivalent formulation of the theory was later given by Holstein and Primakoff [99] based on a transformation that now bears their names, that maps spin operators into bosonic ones. To the lowest order of the theory, this can be regarded as approximating the spin system as a set of harmonic oscillators [100], where spin-wave excitations are captured as bosonic quasiparticles. The spin-wave theory has been highly successful in describing equilibrium ground-states in a wide range of magnetic materials, showing good accordance with experimental data [100]. More recently, the spin-wave theory has been generalized to the time-dependent regime [101, 102], capturing non-equilibrium dynamics in closed quantum systems undergoing unitary time evolution governed by a Hamiltonian and in driven-dissipative systems [97] described by a Lindblad master equation. This work enlarges spin-wave theory to encompass generalized measurements, therefore providing the natural framework for the study of monitored Hamiltonian and Lindbladian long-range systems.
@@
The paper is organized as follows. In Section II we set the stage by defining the class of models we are going to examine and the associated average dynamics, the Lindblad equation. Section III is the core of our work as it contains the essence of the stochastic spin-wave approach. In this Section, we explain how it is constructed in the case of a monitored dynamics described by a quantum state diffusion (QSD). The approach can be transposed in a straightforward way to other types of unravelings. We derive the stochastic equations of motions for the spin-waves whose averaged dynamics is described by the Lindblad equation introduced before. The method is then applied in Section IV where we we study dissipative phase transitions and entanglement phase transitions in long-range spin systems. In order to test the method we use as a benchmark an infinite-range model where numerically exact calculations can be performed. The agreement is excellent. We then proceed with the generic case where our approach allows us to study large systems, well outside the range of a direct simulation. In Section V we discuss how the post-selection problem can be avoided in long-range spin systems by exploiting quantum-classical cross-correlated observables enabled by our spin-wave framework. Finally, we draw our conclusions in Section VI. Some technical details as well as an extension of our method to spin-boson systems are included in the Appendices.
II Monitored long-range spin systems
@@
$\hat{L}_{i}$ are Lindblad operators acting on the $i$-th spin, and $\langle\hat{L}_{i}\rangle\equiv{\rm{Tr}}[\hat{\rho}\hat{L}_{i}]$ denotes the (single-trajectory) expectation value. In Eq. (2), $\hat{H}$ is the Hamiltonian generating the coherent dynamics, whose form we will specify later, and $\mathcal{D}$ denotes the superoperator governing the dissipation, that acts on the density matrix as follows,
@@
We consider the dynamics of a pure state $\hat{\rho}=|\psi\rangle\langle\psi|$. Given the purity preservation of the dynamics (1) (cf. Appendix A), the von Neumann entropy
@@
is a good quantifier of entanglement [103]. Here, $\hat{\rho}_{\frac{N}{2}}\equiv{\rm{Tr}}_{\{1,\ldots,\lfloor N/2\rfloor\}}[|\psi\rangle\langle\psi|]$ denotes the reduced density matrix of half of the system (in a bipartition where one subsystem contains spins indexed from $1$ to $\lfloor N/2\rfloor$).
@@
As shown in the Appendix A, Eq. (1) is an unraveling of the master equation (6) and describes the quantum state diffusion process subjected to a heterodyne-detection monitoring scheme [104, 25].
II.1 Power-law spin model
While the formulation of our spin-wave method will not depend on the particular form of the model, we generally require $\hat{H}$ and $\mathcal{D}$ to be long-range in order for the method to yield an accurate approximation to the exact solution. For concreteness, we study a prototypical driven-dissipative spin model with spatially extended interaction whose strength decays as a power-law according to the distance. The model is defined on a one-dimensional periodic chain with Hamiltonian
$\hat{H}=\omega\hat{S}^{x}+\dfrac{2sJ}{\mathcal{N}}\sum_{i,j}\dfrac{\hat{\sigma}^{z}_{i}\hat{\sigma}^{z}_{j}}{d(i,j)^{\alpha}}\,,$ (7)
where $\omega$ is the amplitude of a collective drive and $J$ is the interaction strength. The factor $2s$ ensures that the mean-field theory (see Appendix B) is $s-$independent, which can be ignored for spin-half ($s=1/2$) systems. Here, we denote total spin operators with the capital $\hat{S}^{\mu}$, $\mu\in\{x,y,z\}$ without the site index subscript
$\hat{S}^{\mu}\equiv\sum_{i}\hat{s}_{i}^{\mu}\,,$ (8)
while the total spin number is denoted by
$S\equiv Ns.$ (9)
The lower-case notation $\hat{s}_{i}^{\mu}$ refers to the spin operator of the $i-$th site, which satisfies the standard $\mathfrak{su}(2)$ algebra
$={\rm{i}}\delta_{ij}\epsilon^{\mu\nu\gamma}\hat{s}^{\gamma}.$ (10)
We denote the normalized spin operators with
$\hat{\sigma}_{i}^{\mu}\equiv\hat{s}^{\mu}_{i}/s\,,$ (11)
which reduce to the standard Pauli matrices in the case of spin-half. The distance on the periodic chain is
$d(i,j)$ $\equiv\min(|i-j|,N-|i-j|)\,,$ (12)
and in the case of $i=j$ we adopt the convention of $d(i,i)=\infty$ such that there is no on-site self-interaction for finite $\alpha$. The Kac normalization [105]
$\mathcal{N}\equiv\dfrac{1}{N}\sum_{i,j}\dfrac{1}{d(i,j)^{\alpha}}\,,$ (13)
ensures a well-defined thermodynamic limit for the interaction Hamiltonian. The power $\alpha$ determines the range of the interaction. In particular, the case of $\alpha=0$ describes an infinite-range model with permutation invariance, and the opposite limit $\alpha\to\infty$ corresponds to an Ising model with nearest-neighbor interaction. The long-range regime corresponds to the case where $10<\alpha<$ 1.
@@ II.1 Power-law spin model
with $S$ serving the role of Kac normalization for the dissipator, $\kappa$ is the dissipation strength, and $\hat{S}^{\pm}\equiv\hat{S}^{x}\pm{\rm{i}}\hat{S}^{y}$ are the collective spin raising and lowering operators. Note that this collective dissipator corresponds to the choice of $f_{ij}=\kappa/S=\mathrm{const}$ and $\hat{L}_{i}=\hat{s}_{i}^{-}=\hat{s}_{i}^{x}-{\rm{i}}\hat{s}_{i}^{y}$ in terms of our generic notation in Eq. (3).
In the mean-field approximation, the system is predicted to exhibit a continuous phase transition from a stationary phase to a time-crystal one. In the stationary phase, the mean-field steady-state $z$-magnetization is
$\dfrac{\langle\hat{S}^{z}\rangle_{\mathrm{MF}}}{S}=-\sqrt{1-\dfrac{\omega^{2}}{16J^{2}+\kappa^{2}}}\,,$ (15)
cf. Appendix B. This stationary phase terminates at a critical point $\omega_{\mathrm{MF}}^{(c)}=\sqrt{16J^{2}+\kappa^{2}}$, beyond which the $z$-magnetization admits only permanently oscillating solutions with zero mean when averaged over long times.
This model has been previously studied in [91] to demonstrate an entanglement phase transition in the unraveled dynamics in the case of $J=0$, which is an infinite-range collective model. In this regime, the system displays an entanglement phase transition coinciding with the dissipative phase transition from the stationary phase to the time-crystal phase. This transition also features a fast saturation of the entanglement entropy, making its experimental detection free from the exponential post-selection overhead. It was also hypothesized in [91] that the above results should hold in the presence of sufficiently long-range interactions, which was supported by numerical evidence for the case $J\neq 0$ and $\alpha\neq 0$ for up to $10$ spin-$1/2$ sites since the exact simulation requires the full $2^{N}$-dimensional Hilbert space and is not feasible for thermodynamically large $N$. In the following section, we develop the theory of spin-wave quantum trajectories, allowing us to overcome the above limitations and tackle the problem in the thermodynamic limit. Our results are presented in Sec. IV.
III Spin-wave theory along quantum trajectories

### v2

@@ Abstract
Measurement-induced phases exhibit unconventional dynamics as emergent collective phenomena, yet their behavior in tailored interacting systems – crucial for quantum technologies – remains less understood. We develop a systematic toolbox to analyze monitored dynamics in long-range interacting systems, relevant to platforms like trapped ions and Rydberg atoms. Our method extends spin-wave theory to general dynamical generators at the quantum trajectory level, enabling access to a broader class of states than approaches based on density matrices. This allows efficient simulation of large-scale interacting spins and captures nonlinear dynamical features such as entanglement and trajectory correlations. We showcase the versatility of our framework by exploring entanglement phase transitions in a monitored spin system with power-law interactions in one and two dimensions, where the entanglement scaling changes from logarithm to volume law as the interaction range shortens, and by dwelling on how our method mitigates experimental post-selection challenges in detecting monitored quantum phases.
@@ I Introduction
Long-range interacting quantum many-body systems have recently been the focus of intense theoretical and experimental activity [1, 2]. Sufficiently non-local interactions between the elementary constituents of a many-body system lead to exotic non-equilibrium phenomena, with compelling signatures in the properties of quantum correlations and entanglement, relaxation dynamics, quantum information scrambling, and ergodicity-breaking properties. Along with their importance in statistical mechanics, long-range quantum systems are central to the rising field of quantum technologies and simulators. Experimental platforms such as trapped ions [3], Bose-Einstein condensates in cavities [4], dipolar [5], polar [6], and Rydberg atoms [7], or driven ultra-cold atomic gases [8] showcase frameworks where the system interactions scale as a power law of the distance.
@@
The dynamics of long-range systems has been investigated in the two limits of unitary and dissipative (Lindbladian) evolutions. In this work, we would like to take a step forward by studying the in-between framework of monitored dynamics. Here, the system evolution is interspersed with quantum measurements, whose outcomes are stochastic. As a result, the system is described by a quantum trajectory conditional to the measurement registry. Averaging the state over the trajectory ensemble recasts a dissipative dynamics, therefore presenting a convenient framework for the study of physical observables in Lindblad dynamics [9, 10, 11]. It was more recently realized, in the study of many-body systems, that the quantum trajectory ensemble contains richer information than the mean state, showcased by the several collective phenomena encoded in beyond-average statistical features. The cornerstone examples are monitored quantum phases and measurement-induced phase transitions [12, 13, 14], characterized by non-linear functions of the trajectories, such as entanglement probes or trajectory correlation functions. Intensive work gathered insights on the monitored phases in local random quantum circuits [15, 16, 17, 18, 19, 20] and non-interacting systems [21, 22, 23], while few works have addressed monitored dynamics in long-range interacting systems. Refs. [24, 25] studied Clifford circuits with two-qubit gates entangling distant sites with a probability of decaying as a power-law of their distance. The Floquet dynamics of interacting spin systems with measurements and feedback was considered in a model with entanglement and dissipative phase transitions [26], while monitored long-range free fermions have been studied in Refs. [27, 28]. All these results demonstrate that long-range couplings strongly affect the monitored phases and their transitions: they are relevant in the renormalization group sense.
@@
Experimentally detecting monitored phases, on the other hand, is challenging, and so far limited to few pioneering works on trapped-ions [29] and superconducting platforms [30, 31]. This limitation has a fundamental origin known as the post-selection problem. Observing beyond-average moments of the quantum trajectories requires reproducing the same sequence of measurement outcomes multiple times to collect enough statistics and obtain trajectory-average values. This task is formidable, as the probability of reproducing the same trajectory is exponentially small in system size and time scale.
@@
The post-selection barrier is in general ineludible for the study of generic many-body monitored systems, and the quest for methods that can mitigate it is an active research line. For instance, the presence of feedback dynamics may be designed to imprint the measurement-induced transition into the density matrix [32, 33], albeit in general leads to separate types of phase transitions [34, 35, 36, 37, 38, 39, 40]. A complementary path is available for systems that can be classically simulated. When the post-processing is efficient and faithful, it allows to elude the post-selection [41, 42, 43, 44, 45, 46, 47]. There are special cases in which the system itself is designed to be immune to post-selection. In [48] some of us showed that there is a class of infinite-range spin systems where monitored many-body dynamics can be efficiently realized with a post-selection overhead scaling, at most, as a power of the system size. It was argued that this fortunate case was not specific to that model, but it applies to a broad class of monitored systems with an underlying semi-classical dynamics.
@@
As long-range systems are many-body and interacting by construction, their theoretical study also poses a formidable challenge by itself. One may resort to approximative numerical methods, such as those based on matrix-product states [49, 50, 51] and other types of tensor networks [52], representing the state-of-the-art for one-dimensional systems and in low entropy scenarios. These methods, while being “generic”, struggle in capturing correlations built up in nonequilibrium dynamics (for example in a phase transition) and generally suffer from poor complexity scaling in two or higher spatial dimensions. The method we propose, on the other hand, exploits the simplifications granted by long-range interactions: they typically lead to underlying semi-classical dynamics, allowing quantum excitations and fluctuations to be treated perturbatively on top of a classical model. These observations lead to a class of approximation methods known as the spin-wave theory. Historically, the spin-wave theory was first introduced by Bloch [53] in 1932 for ferromagnetic spin systems in equilibrium. An equivalent formulation of the theory was later given by Holstein and Primakoff [54]. The spin-wave theory has been highly successful in describing equilibrium ground states in a wide range of magnetic materials, showing good accordance with experimental data [55]. More recently, it has been generalized to the time-dependent regime [56, 57], capturing non-equilibrium dynamics in closed quantum systems undergoing unitary time evolution governed by a Hamiltonian and in driven-dissipative systems [58] described by a Lindblad master equation.
This work enlarges spin-wave theory to encompass generalized measurements, therefore providing a natural framework for the study of monitored Hamiltonian and Lindbladian long-range systems. We demonstrate the method on both one- and two- dimensional spin lattices and on a spin-boson model, exploring different monitored phases where the entanglement scaling ranges from logarithmic to volume law. Moreover, as our method is designed on the level of single quantum trajectories, it also provides a much more accurate representation of the Lindblad dynamics, allowing to capture non-Gaussian corrections that are by default lost in semiclassical approaches approximating directly on the level of the density matrix [58], offering a broad spectrum of applications.
@@
The paper is organized as follows. In Section II.1 we set the stage by defining the class of models we are going to examine and the associated Lindblad equation. Section II.2 is the core of our work as it contains the essence of the stochastic spin-wave approach. In this Section, we explain how it is constructed in the case of a monitored dynamics described by a quantum state diffusion. The method is then applied in Section II.3 where we benchmark it with exactly solvable cases and we apply it to the study of dissipative phase transitions and entanglement phase transitions in long-range spin systems. In Section II.4 we discuss how the post-selection problem can be avoided in long-range spin systems by exploiting quantum-classical cross-correlated observables enabled by our spin-wave framework. Finally, we draw our conclusions in Section III. Some technical details are included in the Methods section.
Figure 1: Sketch of a long-range spin model [panel (a)] and the representation of its monitored dynamics in the spin-wave approximations [panels (b) and (c)]. (a) The spins on a $d$-dimensional lattice are collectively driven (at amplitude $\omega$) and interact via $z-z$ coupling, whose strength decays as a power law of the distance (with exponent $\alpha$). The system is subjected to a collective dissipation (at rate $\kappa$) and continuously monitored. In the infinite-range or long-range regime ($0\leq\alpha\leq d$), the state of the system can be represented in the spin-wave approximations [panel (b)], where spin fluctuations around the collective polarization are bosonized. Here, $Oxyz$ depicts the lab frame and $O{\tilde{x}}{\tilde{y}}{\tilde{z}}$ is the instantaneous frame where ${\tilde{z}}$ aligns with the collective spin polarization. Panel (c) represents the evolution of the state under both the Hamiltonian dynamics and the continuous monitoring. In the proposed method, the angles $(\theta,\phi)$ of the instantaneous frame $O{\tilde{x}}{\tilde{y}}{\tilde{z}}$ are updated to match the collective spin direction after its evolution in each infinitesimal time step. The dashed (solid) straight arrow represents the ${\tilde{z}}$ axis before (after) the re-alignment step and the curved arrows represent the paths $\gamma_{1}$ and $\gamma_{2}$ as defined in Eq. (25) for evaluating the integral of Eq. (24).
II Results
II.1 Monitored long-range spin systems
@@
$\hat{L}_{i}$ is the Lindblad operator acting on the $i$-th spin, and $\langle\hat{L}_{i}\rangle\equiv{\rm{Tr}}[\hat{\rho}\hat{L}_{i}]$ denotes the (single-trajectory) expectation value. In Eq. (2), $\hat{H}$ is the Hamiltonian generating the coherent dynamics, whose form will be specified later, and $\mathcal{D}$ denotes the superoperator governing the dissipation, that acts on the density matrix as follows,
@@
We consider the dynamics of a pure state $\hat{\rho}=|\psi\rangle\langle\psi|$. Given the purity preservation of the dynamics (1) (cf. Supplementary Note A), the von Neumann entropy
@@
is a good quantifier of entanglement [59]. Here, $\hat{\rho}_{\frac{N}{2}}\equiv{\rm{Tr}}_{\{1,\ldots,\lfloor N/2\rfloor\}}[|\psi\rangle\langle\psi|]$ denotes the half-system reduced density matrix (in a bipartition where one subsystem contains spins indexed from $1$ to $\lfloor N/2\rfloor$).
@@
As shown in the Supplementary Note A, Eq. (1) is an unraveling of the master equation (6) and describes the quantum state diffusion process subjected to a heterodyne-detection monitoring scheme [60, 10].
II.1.1 Power-law spin model
While the formulation of our spin-wave method will not depend on the particular form of the model, we generally require $\hat{H}$ and $\mathcal{D}$ to be long-range in order for the method to yield an accurate approximation to the exact solution. For concreteness, we study a prototypical driven-dissipative spin model with spatially extended interaction whose strength decays as a power-law according to the distance, as sketched in Fig. 1 (a). The model is defined on a $d$-dimensional periodic lattice of $N=L^{d}$ spins at positions $\bm{r}_{i}\equiv(r_{i}^{(1)},\ldots,r_{i}^{(d)})$ where $r_{i}^{(p)}=1,\ldots,L$. The Hamiltonian is
$\hat{H}=\omega\hat{S}^{x}+\dfrac{2sJ}{\mathcal{N}}\sum_{i,j}\dfrac{\hat{\sigma}^{z}_{i}\hat{\sigma}^{z}_{j}}{{\lVert\bm{r}_{i}-\bm{r}_{j}\rVert}^{\alpha}}\,,$ (7)
where $\omega$ is the amplitude of a collective drive and $J$ is the interaction strength. The factor $2s$ ensures that the mean-field theory (see Supplementary Note B) is $s-$independent, which can be ignored for spin-half ($s=1/2$) systems. Here, we denote total spin operators with the capital $\hat{S}^{\mu}$, $\mu\in\{x,y,z\}$ without the site index subscript $\hat{S}^{\mu}\equiv\sum_{i}\hat{s}_{i}^{\mu}\,$, while the total spin number is denoted by $S\equiv Ns$. The lower-case notation $\hat{s}_{i}^{\mu}$ refers to the spin operator of the $i-$th site, which satisfies the standard $\mathfrak{su}(2)$ algebra $[\hat{s}_{i}^{\mu},\hat{s}_{j}^{\nu}]={\rm{i}}\delta_{ij}\epsilon^{\mu\nu\gamma}\hat{s}^{\gamma}$. We denote the normalized spin operators with $\hat{\sigma}_{i}^{\mu}\equiv\hat{s}^{\mu}_{i}/s\,$, which reduce to the standard Pauli matrices in the case of spin-half. The distance on the periodic lattice is $\lVert\bm{r}_{i}-\bm{r}_{j}\rVert\equiv\sqrt{\sum_{p=1}^{d}\min(|r_{i}^{(p)}-r_{j}^{(p)}|,L-|r_{i}^{(p)}-r_{j}^{(p)}|)^{2}}$, and in the case of $i=j$ we adopt the convention of ${\lVert\bm{r}_{i}-\bm{r}_{i}\rVert}=\infty$ such that there is no on-site self-interaction for finite $\alpha$. The Kac normalization [61] $\mathcal{N}\equiv\frac{1}{N}\sum_{i,j}{\lVert\bm{r}_{i}-\bm{r}_{j}\rVert}^{-\alpha}\,$, ensures a well-defined thermodynamic limit for the interaction Hamiltonian. The power $\alpha$ determines the range of the interaction. In particular, the case of $\alpha=0$ describes an infinite-range model with permutation invariance, and the opposite limit $\alpha\to\infty$ corresponds to an Ising model with nearest-neighbor interaction. The long-range regime corresponds to the case where $0<\alpha\leq d$ [62].
@@ II.1.1 Power-law spin model
with $S$ serving the role of Kac normalization for the dissipator, $\kappa$ is the dissipation strength, and $\hat{S}^{\pm}\equiv\hat{S}^{x}\pm{\rm{i}}\hat{S}^{y}$ are the collective spin raising and lowering operators. Note that this collective dissipator corresponds to the choice of $f_{ij}=\kappa/S=\mathrm{const}$ and $\hat{L}_{i}=\hat{s}_{i}^{-}=\hat{s}_{i}^{x}-{\rm{i}}\hat{s}_{i}^{y}$ in terms of our generic notation in Eq. (3), which also fixes the stochastic unraveling via Eq. (1).
The quantum state of the spin system lives in a $2^{N}$-dimensional Hilbert space, making the exact solution intractable in practice for thermodynamically large $N$. In the following section, we elaborate the theory of spin-wave quantum trajectories, allowing us to overcome the above limitations and tackle the problem in the thermodynamic limit.
II.2 Spin-wave theory along quantum trajectories